OpenStudy (anonymous):

simplify the expression. 14ab^3 -------- 7a^-2b^-1

OpenStudy (e.mccormick):

Do you know the rules for negative exponents?

OpenStudy (anonymous):

no.

OpenStudy (e.mccormick):

\[\frac{1}{x^{-1}}=x\text{ and }x^{-1}=\frac{1}{x}\]

OpenStudy (e.mccormick):

So on yours:\[\frac{14ab^3}{7a^{-2}b^{-1}}\]Do you see how the negative exponents will become positive if they are moved to the top?

OpenStudy (anonymous):

oh yes! because of the negative exponent, i get what you're saying now.

OpenStudy (anonymous):

would you mind explaining it step by step?

OpenStudy (e.mccormick):

When they are on the top, then you do combined like terms. The only other thing to do would be to simplify the 14 over 7 part.

OpenStudy (anonymous):

so would it be \[2a ^{-2}b ^{-3}\] ?

OpenStudy (e.mccormick):

Not quite... The 2 is great! But you did not do the exponents correctly.

OpenStudy (anonymous):

\[2a ^{3}b ^{4}\] ?

OpenStudy (e.mccormick):

Yes!

OpenStudy (e.mccormick):

That is the right way to do it!

OpenStudy (anonymous):

thank you!

OpenStudy (e.mccormick):

So just to review it:\[\frac{14ab^3}{7a^{-2}b^{-1}}\implies\\ \, \\ \frac{2ab^3}{a^{-2}b^{-1}}\implies \\ \, \\ \frac{2aa^2b^3b}{1}\implies\\\, \\ 2a^3b^4\]

OpenStudy (anonymous):

thank you so much, i really appreciate it! (:

OpenStudy (e.mccormick):

NP. have fun!